OpenStudy (anonymous):

x^2/16+ y^2/36 =1 Vertices:______ Co-Vertices:______ Foci:_______

OpenStudy (anonymous):

\[\frac{ x^2 }{ 16 } + \frac{ y^2 }{ 36 } = 1\]

OpenStudy (anonymous):

vertices:(0,6), (0,-6) co vertices: Foci:(0, -2sqrt(5)) , (0, 2sqrt(5))

OpenStudy (anonymous):

i dont think there is a covertices

OpenStudy (anonymous):

Thanks again!